Ask
your own question, for FREE!
Mathematics
16 Online
@Michele_Laino A box holds 42 pieces of red construction paper and 18 pieces of blue construction paper. A piece of paper is pulled out of the box at random and then placed back in the box. How many times can you expect that a piece of red paper will be pulled out of the box in 110 tries? A. 48 B. 55 C. 60 D. 77
Still Need Help?
Join the QuestionCove community and study together with friends!
I got D
hint: what is the probability to get a red paper?
77?
42/(42+18)=7/10 7/10*110=77 red papers
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
13 hours ago
0 Replies
0 Medals
4 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
2 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal